# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v2

@@ Abstract
In this paper, we explore constructions for codes that correct two bursts of deletions, with each burst having length exactly $b$. Previously, the best known construction, derived using the syndrome compression technique, achieved a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits. In this work, we present new constructions for all $q\geq 2$ that achieve redundancy at most $7\log n+O(\log\log n)$ when $b>1$. Additionally, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@ I Introduction
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. To the best of our knowledge, there is no explicit results about such codes. A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. In this paper, we construct codes which outperform the one derived by the syndrome compression technique. Our contributions include:
  • A binary code of length $n$ with redundancy at most $7\log n+(15b-5)\log\log n+O(1)$;
  • A $q$-ary code of length $n$ with redundancy at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $7\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
@@ II Preliminaries
    the two bursts happen sequentially: from $\mathbf{x}$, we get $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then we get $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$ from $\mathbf{z}$;
@@
    the two bursts happen simultaneously: from $\mathbf{x}$, we get $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n]. This definition is motivated by the nanopore sequencing technology, in which a DNA strand pass a nanoscale hole nucleotide by nucleotide [36, 37]. So we can assume that the two bursts do not overlap.
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that see that (D1) covers (D2). Next, we show that (D2) also covers (D1).
@@ 1. $(\textup{{P}}1)$
Before proceeding to subsequent sections, we introduce a well-known result from number theory, which will be frequently used.
@@ III Bounds
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [39, page 100, Theorem 1] that
@@
In [40, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
@@
Now it follows from Equations 4 and 5 that
@@
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 3 and 6. ∎
@@ 2. $(2)$
For the first step, for all $1\leq j<r\left(A(\mathbf{x})_{1}\right)$, define
@@
which can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{b-1}$. For $s=0,1$, let
$K_{s}=\left\{2\leq j\leq r\left(A(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime. For $s=0,1$, Define
$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}f_{j-1}(\mathbf{x})\pmod{2N_{1}},$ (8)
$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)f_{j-1}(\mathbf{x})\pmod{Q_{1}}.$
In the following, we may assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4.

### v3

@@ Abstract
In this paper, we study codes that correct two bursts of deletions, with each burst having length exactly $b$, where $b>1$. Previously, the best known construction, derived using the syndrome compression technique, achieved at most $7\log n+O\left(\log n/\log\log n\right)$ redundant bits. In this work, we develop new idea to construct $q$-ary codes that achieve redundancy at most $6\log n+O(\log\log n)$ for all $b>1$ and $q\geq 2$. At last, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@ I Introduction
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. To the best of our knowledge, there is no explicit results about such codes. A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] gave a binary code with redundancy at most $7\log n+O(1)$ and [19] gave a binary code with redundancy at most $4\log n+O(\log\log n)$. Note that in a burst of deletions, all deletions occur consecutively. Therefore, intuitively, it is reasonable to deem that there is no big difference between two-deletion correcting codes and codes for correcting two $b$-burst-deletions (where $b>1$). This raises a natural question: for $b>1$, is there a construction which leads to codes which are as good as or even better than the one given by the syndrome compression technique. Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
  • A binary code of length $n$ with redundancy at most $6\log n+(15b-5)\log\log n+O(1)$;
  • A $q$-ary code of length $n$ with redundancy at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
@@ II Preliminaries
    the two bursts are caused by two channels: $\mathbf{x}$ passes the first channel, resulting in $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then $\mathbf{z}$ passes the second channel, resulting in $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$;
@@
    the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
Remark II.1
@@ 1. $(\textup{{P}}1)$
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory, which will be frequently used. The second one will be used in Sections IV-A and V-A.
@@
Let $N>2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Sections IV-A and V-A, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
Lemma II.7
Let $Q\geq N$ be the smallest prime. For all $0\leq a_{1},a_{2},b_{1},b_{2}<Q$ and $0\leq c<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
  • (C1)
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\equiv a_{1}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\equiv a_{2}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\equiv b_{1}\pmod{Q}$,
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\equiv b_{2}\pmod{Q}$;
  • (C2)
    $W(\mathbf{x})\equiv c\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+2N\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts of erasures. In particular, there is a function $\mathcal{E}_{1}:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+6\log N+4}$, computable in linear time, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\mathcal{E}_{1}(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
Proof:
Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
$\mathcal{E}_{1}(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\pmod{Q},\right.$
$\left.\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\pmod{Q},W(\mathbf{x})\pmod{nN^{2}}\right).$
Here, we view $\mathcal{E}(\mathbf{x})$ as a binary vector. By Lemma II.6, we have $N\leq Q<2N$. Now it is easy to see that the length of $\mathcal{E}(\mathbf{x})$ is at most $\log n+6\log N+4$. It remains to prove the correctness of $\mathcal{C}$.
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Suppose the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$. By definition, the two bursts in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$.
To decode $\mathbf{x}$ from $\mathbf{y}$. We first calculate the two erased symbols in each row. Notice that $N>2$ and hence $Q$ is an odd prime. Suppose $\sigma_{1},\sigma_{2}\in[0,N-1]$ satisfy
$\sigma_{1}+\sigma_{2}\equiv t_{1}\pmod{Q},$ (1)
$\sigma_{1}^{2}+\sigma_{2}^{2}\equiv t_{2}\pmod{Q}$
where $t_{1}$ and $t_{2}$ are given. Then we can uniquely recover the (multi)set $\{\sigma_{1},\sigma_{2}\}$ by solving system (1). Now we can determine the erased symbols in $A(\mathbf{y})_{1}$ and $A(\mathbf{y})_{2}$ by (C1). Suppose that the two erased symbols in $A(\mathbf{y})_{1}$ are $\alpha_{1},\alpha_{2}$, and the two erased symbols in $A(\mathbf{y})_{2}$ are $\alpha_{3},\alpha_{3}$. It remains to determine positions of $\alpha_{1}$ and $\alpha_{2}$ in $A(\mathbf{x})_{1}$, and positions of $\alpha_{3}$ and $\alpha_{4}$ in $A(\mathbf{x})_{2}$. We can assume $\alpha_{1}\neq\alpha_{2}$ or $\alpha_{3}\neq\alpha_{4}$, since otherwise, we can uniquely recover $\mathbf{x}$.
There are at most four choices for the positions of $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and for each choice, we get a candidate for $\mathbf{x}$. We need to show that among these candidates, there is exactly one equals $\mathbf{x}$. Firstly, by the relationship of $\mathbf{x}$ and $\mathbf{y}$, it is clear that there is at least candidate equals $\mathbf{x}$. Next, we will show there is at most one such candidate. To that end, let $(\beta_{1},\beta_{2})$ be a permutation of $(\alpha_{1},\alpha_{2})$, and $(\beta_{3},\beta_{4})$ be a permutation of $(\alpha_{3},\alpha_{4})$, where $(\beta_{1},\beta_{2})\neq(\alpha_{1},\alpha_{2})$ or $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Let $\mathbf{u}$ be obtained from $\mathbf{y}$ by assigning $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. Let $\mathbf{v}$ be obtained from $\mathbf{y}$ by assigning $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. By assumption, we have $\mathbf{u}\neq\mathbf{v}$. We need to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ is not a codeword in $\mathcal{C}$. Let $\Delta=W(\mathbf{u})-W(\mathbf{v})$. Then we have
$\Delta=i_{1}(\alpha_{1}-\beta_{1})+i_{2}(\alpha_{2}-\beta_{2})+2N\left(i_{3}(\alpha_{3}-\beta_{3})+i_{4}(\alpha_{4}-\beta_{4})\right).$
To complete the proof, it is sufficient to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ does not satisfies (C2). Suppose on the contrary that $W(\mathbf{u})\equiv c\equiv W(\mathbf{v})\pmod{nN^{2}}$. Then we have $\Delta\equiv 0\pmod{nN^{2}}$.
If $(\beta_{1},\beta_{2})=(\alpha_{1},\alpha_{2})$, it holds that $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Therefore, we conclude that $\alpha_{3}\neq\alpha_{4}$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then it follow that $\left|\Delta\right|=2N(i_{4}-i_{3})\left|\alpha_{3}-\alpha_{4}\right|$ and hence $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)<nN^{2}$, which is a contradiction. Similarly, if $(\beta_{3},\beta_{4})=(\alpha_{3},\alpha_{4})$, we have $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)<nN^{2}$, which is a contradiction.
Now suppose $\alpha_{1}\neq\alpha_{2}$, $\alpha_{3}\neq\alpha_{4}$, $(\beta_{1},\beta_{2})=(\alpha_{2},\alpha_{1})$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then we have $\Delta=(i_{2}-i_{1})(\alpha_{2}-\alpha_{1})+2N(i_{4}-i_{3})(\alpha_{4}-\alpha_{3})$. Without loss of generality, suppose $\alpha_{1}<\alpha_{2}$. If $\alpha_{3}<\alpha_{4}$, then $0<\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)+\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)\leq n(N^{2}-1)$, which is a contradiction. Now suppose $\alpha_{3}>\alpha_{4}$. Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to verify that $0<-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(2N(N-1)-1)+2N(N-1)<nN^{2}$, which is a contradiction. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, we have $i_{2}\geq i_{1}+2$. Then it follows that
$0<2N(\alpha_{3}-\alpha_{4})-2(\alpha_{2}-\alpha_{1})\leq-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor\left(2N(N-1)-1\right)-2N<nN^{2},$
which is a contradiction. ∎
@@ III Bounds
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [37, page 100, Theorem 1] that
@@
In [38, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
where (a) follows from the well-known result (see the proof of [40, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [41, Proposition 3.3.3] that
@@
Now it follows from Equations 5 and 6 that
@@
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 4 and 7. ∎
@@ 2. $(2)$
For the first step, for all $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
@@
which can be viewed as an integer in $\left[0,N-1\right]$, where $N=2^{b-1}$. For uniform discussion, we also let $f_{0}(\mathbf{x})=0$. Now define $L(\mathbf{x})=f_{0}(\mathbf{x})f_{1}(\mathbf{x})\cdots f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})\in[0,N-1]^{r\left(A(\mathbf{x})_{1}\right)+1}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\ldots,f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})$. To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$. By Equation 8, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
In the following, we always assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4. Now we have $j_{2}-1\geq j_{1}-1+2$. Then it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(A(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+1)+6\log N+4}=2^{\log n+O_{b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.